Given is a historical time series for iob services demand in the prior 6 months. Month Demand 698 815 716 613 657 865 4 The Period 4 forecast by using exponential smoothing with 0.25 alpha (in 2 decimal places)
Answer #1

F2 = 0.25*698 + 0.75*698

= 698

F3 = 0.25*815 + 0.75*698

= 203.75 + 523.5

= 727.25

F4 = 0.25*716 + 0.75*727.25

= 179 + 545.4375

= 724.4375